# CRAFT Your Dataset: Task-Specific Synthetic Dataset Generation Through Corpus Retrieval and Augmentation

### v1

@@ 4.2.1 Metrics
QA Tasks:
All multiple-choice QA tasks are evaluated using accuracy. We follow the evaluation approach of MMLU and assess the logarithmic probabilities for the vocabulary options corresponding to the letter labels of the answer choices (Hendrycks et al., 2021). Accordingly, greedy decoding without temperature scaling is performed. Depending on the number of answer choices, this evaluation can range from options A and B to options A through E.
Generative Tasks:
Automated metrics like ROUGE (Lin, 2004) and METEOR (Banerjee and Lavie, 2005) are resource-efficient as evaluation metrics but face limitations in generative tasks. They rely heavily on n-gram overlap, which may not accurately reflect the true quality of the generated text (Barbella et al., 2021). They also assume that the reference text provides a complete and accurate representation of the desired output, which is not always guaranteed (Graham, 2015; Sai et al., 2019). Reference texts can be low-quality, contain errors and ambiguities, which leads to unreliable evaluation results. This issue is illustrated by the findings of Sottana et al. (2023), who demonstrated that human reviewers frequently rate gold-standard benchmark answers among the worst answer options. This issue worsens when generated texts differ significantly in length from references, as subsequence-based metrics struggle to capture such variations (Celikyilmaz et al., 2020).
As an alternative, we opt to evaluate generations using LLMs as a judge (Eldan and Li, 2023). In this setup, the LLM effectively acts as a human annotator, providing a binary preference score for each pair of outputs, resulting in a win rate as the final metric (Chiang et al., 2024). This approach has become increasingly common: LLMs have been successfully employed as annotators while demonstrating high inter-rater reliability (Hackl et al., 2023; Sottana et al., 2023; Liu et al., 2023).
For general-purpose outputs, we use the popular Alpaca-Eval benchmark (Li et al., 2023) that evaluates multiple LLMs on about 650 human-curated questions (Dubois et al., 2023). We select Llama 3 70B (Dubey et al., 2024) as our annotator model due to its open nature and cost-efficiency for high-volume experiments. As of July 2024, Llama 3 70B ranks 4th in human agreement with a score of 67.5, close to customized GPT-4 versions at 69.2.
@@ 4.2.2 Datasets
To assess the quality of our synthetically generated datasets, we compare the performance of models trained on them to those trained on human-curated datasets.
BioQA:
We use the 800 sample test split from the biology subsection of ScienceQA (Lu et al., 2022). ScienceQA sourced expert-curated question-answer pairs from online learning platforms, ensuring a high level of quality and accuracy. The dataset’s answer options range from two to five, have a single correct answer per question, and are randomized to prevent pattern recognition.
MedQA:
We use the 4183 samples from the MedMCQA validation split (Pal et al., 2022). The dataset is comprised of entrance exam questions to two of India’s postgraduate institutions. All dataset samples are sourced from preparation or real exams created by medical professionals. All questions have 4 answer options with one correct answer option per question.
CSQA:
We benchmark on the 2541 validation set samples in CommonsenseQA 2.0 (Talmor et al., 2021). The dataset was generated through a gamified but controlled question generation process, where players were incentivized to design challenging yes/no questions by earning points when beating an AI model. The generated questions were then cross-validated by other players, independent validators, and another model, to ensure that the questions were well-formed, answerable, and representative of common sense.
RecipeGen:
We randomly sample 1000 samples from the higher-quality subsection of the RecipeNLG dataset (Bień et al., 2020). The recipes were scraped from cooking websites and post-processed using a fine-grained and standardized cleaning and formatting process to ensure correctness. Each recipe features a title, the list of required ingredients, as well as the steps to produce the meal. We only include samples not present in C4, based on the URL in each dataset.
Figure 3: Performance scaling with increasing data size across multiple tasks when running CRAFT with 8 (solid blue line) and 32 (dashed black line) few-shots. The graphs illustrate consistent improvement as the training data size increases from few-shots (XS) to 25,000 synthetic task samples (XL). CRAFT-trained models consistently outperform or match the Mistral 7B Instruct v0.2 baseline (dotted red line). Shaded regions represent the standard deviation across three runs.
@@
We use 1000 samples from the test split of the CNN-DailyMail dataset (See et al., 2017). The dataset is commonly used for summarization because it consists of articles and stories from CNN and DailyMail alongside their highlights in abstract or bullet point format presented at the top of newspaper pages or websites.
5 Results
5.1 Baselines
We compare the performance of CRAFT, trained on our synthetic datasets, against three baselines. The few-shot baseline is a model fine-tuned only on the XS size CRAFT dataset, with human-curated few-shots. It serves as the primary baseline since this model uses all human-curated data available in our pipeline.
The second baseline is the instruction-tuned model, Mistral 7B Instruct v0.2 (Jiang et al., 2023), which has been fine-tuned on proprietary instruction-following datasets that mix various tasks and sources. This baseline provides a meaningful comparison, as it is similar in size and instruction-tuned like CRAFT models, though it is trained on undisclosed datasets of unknown quality and quantity. Thus, matching or exceeding the performance of instruction-tuned models with our synthetic data would indicate that CRAFT can produce high-quality datasets.
The upper bound of expected performance is fine-tuning the models on the in-domain human-curated training splits from the chosen evaluation datasets. This baseline represents the optimal performance achievable with human-quality datasets.
5.2 Scaling the Data
We report the performance gain when scaling up our training data in Figure 3. We report the mean and standard deviation across three seeds. We observe consistent improvements across four tasks as we increase the data size. Relative to the baseline models trained with only few-shot examples, we see improvements of 17% (from 66.9 to 78.1), 12% (from 55.3 to 62.1), 23% (from 39.1 to 48.0), and 124% (from 43.7 to 97.9) for BioQA, CSQA, MedQA, and Summarization, respectively. This shows that CRAFT can be used for diverse tasks, starting with just a few curated examples. We also find appropriate scaling for each set of examples, ranging from 100 to 25,000 across all tasks. Additionally, that models trained with fewer examples (32, 100) exhibit much more variance than those trained with 5,000 and 25,000 examples, as indicated by the gray regions in the plots that visualize the standard deviation.
For all tasks, we achieve results that are are clearly better of comparable to Mistral Instruct. It is worth noting that CRAFT uses an LLM in a limited way (to restructure and rewrite existing corpora) that seem to exclude the possibility that distillation may have played a role here. However, even if distillation were to be considered the reason for good CRAFT performance, the results indicate otherwise: we use the same model (Mistral Instruct) to paraphrase existing corpora examples but achieve even stronger results. Finally, we observe that CRAFT models outperform those trained with official human-curated large-scale data in Summarization. For other tasks, while we observe lower performance than with official data, we speculate that this could be due to in-domain evaluation for official human-curated data. We use their test split to evaluate our models, which may give these models an unfair advantage. We investigate this further in the next section.
Dataset Baseline CRAFTXL
In-domain 89.9 78.1
MMLUMedical Genetics 60.0 69.0
MMLUAnatomy 55.6 57.0
MMLUHigh School Biology 69.3 67.4
MMLUCollege Biology 66.7 74.3
MMLU-Avg 62.9 66.9
Table 1: OOD evaluation of baseline and CRAFTXL.
5.3 OOD Generalization and Data Contamination
We now report experiments to understand the level of data contamination or similarity between test and training examples in the experiments introduced in Figure 3. We conduct 5-gram weighted Jaccard similarity analyses between CRAFT datasets and the test data. For each sample, we combine the instruction and output and gather 5-gram frequencies for the whole dataset. We then calculate the Jaccard similarity between the 5-gram frequency distributions of the respective CRAFT and test dataset, where 5-grams receive weight proportional to their frequency.
This analysis shows that all CRAFT datasets have less than 0.4% similarity with the task test set, whereas the original human-authored datasets show much higher similarities: BioQA (17.9%); CSQA (4.4%), MedQA (1.1%), and Summarization (0.3%); this may indicate some overlap between train and test splits. Since we have curated the few-shots manually, rather than copying from existing datasets, our low overlap (0.4%) may be expected. To further investigate CRAFT’s performance improvement, we select four out-of-domain datasets for the biology question answering task and compare CRAFT with the human-curated baseline.
In Table 1, we compare the human-curated baseline with CRAFT on the in-domain dataset for the baseline and four out-of-domain (OOD) datasets selected from MMLU Hendrycks et al. (2021). Although the baseline outperforms CRAFT by more than 11% in the biology subset of the ScienceQA test set (in-domain for the baseline), CRAFT outperforms the baseline on 3 out of 4 OOD datasets. On average, CRAFT outperforms the baseline by 4%.
Task Model CRAFTXS CRAFTS CRAFTM CRAFTL CRAFTXL
BioQA Mistral 66.88±1.24 65.71±1.74 (-1.74%) 71.00±1.87 (6.17%) 76.75±0.64 (14.76%) 78.13±0.35 (16.82%)
Llama 64.38±1.98 69.09±0.78 (7.32%) 70.38±0.57 (9.32%) 73.38±1.24 (13.99%) 76.63±1.22 (19.03%)
CSQA Mistral 55.33±5.09 56.47±4.76 (2.06%) 60.68±0.31 (9.67%) 61.77±0.80 (11.63%) 62.10±0.54 (12.23%)
Llama 56.72±0.82 54.60±0.44 (-3.74%) 59.57±0.85 (5.02%) 60.45±0.40 (6.58%) 61.64±0.64 (8.67%)
MedQA Mistral 39.17±0.48 41.81±1.79 (6.72%) 46.49±1.13 (18.68%) 46.16±0.27 (17.83%) 47.99±0.36 (22.52%)
Llama 49.65±0.35 49.70±0.07 (0.10%) 52.38±0.11 (5.49%) 52.07±0.42 (4.86%) 52.82±0.29 (6.37%)
Summ Mistral 43.73±2.78 52.80±3.35 (20.73%) 91.90±1.02 (110.14%) 96.23±0.40 (120.05%) 97.90±0.37 (123.86%)
Llama 31.70±3.97 28.20±3.28 (-11.04%) 88.27±1.18 (178.44%) 96.63±0.24 (204.84%) 97.17±0.38 (206.52%)
Table 2: Performance comparison of Llama and Mistral across different tasks and few-shot settings. Values show accuracy ± standard deviation, with relative improvement in parentheses compared to the CRAFTXS model.
5.4 Negative Results: Recipe Generation
Out of the five tasks we selected, we observe non-scaling behavior in one task: Recipe Generation. While our manually curated few-shots are of high quality, we see a drop when scaling from 32 to 25,000 examples, as illustrated in Figure 4. CRAFT’s performance is still better than the baseline with official human data, which means that the final dataset is usable. However, we explore why this reverse scaling occurs and examine the drop in performance.
Figure 4: Non-scaling behavior observed in the RecipeGen task. Unlike the general trend, performance does not consistently improve with increasing data size. We report the overall dataset quality with a solid orange square line. The graph shows an inverse relationship between dataset size and quality. This anomaly in scaling behavior is attributed to a decline in data quality as the number of examples increases.
An initial analysis suggested that the CRAFT pipeline tends to find less relevant examples over time. We conducted automated data quality analysis to analyze this on a larger scale. For 500 randomly sampled instructions from different sizes of CRAFT datasets (i.e., the training sets), we used the Llama 3 8B Instruct model to answer the instructions. Then, using Llama 3 70B Instruct as a judge, we compared win rates, i.e., which output the model preferred: the gold output in the CRAFT dataset or the output generated by Llama 3 8B Instruct. We report the average win rate against the Llama outputs as the data quality metric. Higher scores indicate that the pipeline created higher quality output than Llama 3 8B Instruct’s answers.
We observe that data quality drops when scaling up to 25,000 examples. The 100 and 500 example sets have win rates around 0.4, while the win rate for 25K drops to 0.3. We believe this is the cause of the performance drop with scaling. While the final dataset is still useful (it outperforms the baseline with official human data), the next version of CRAFT should include either effective stopping criteria or additional validators for quality.
5.5 Base Model Comparison
In previous sections, we fine-tuned CRAFT models using the pretrained Mistral 7B model. Now, we repeat the experiments using the pretrained Llama 3 8B model. We observe similar trends across all tasks, and the relative improvement is comparable when scaling up from few-shots to 25,000 examples, as illustrated in Table 2.
In all experiments, we manually curated 32-shot examples and expanded our synthetic data examples from that point. However, even curating 32 examples can be time-consuming. We can limit the number of few-shots to just eight examples to bootstrap the CRAFT process. Figure 3 shows the results compared to the CRAFT pipeline with 32-shots. While the final results with 25,000 examples are slightly lower for 8-shots, the trend is similar for both 8- and 32-shot examples. We observe that the model trained with 25,000 samples, based on running CRAFT with 8 few-shots, significantly outperforms the model trained with only 32 few-shots (i.e., no extra synthetic data). This suggests that if there are time and resource limitations, using CRAFT with fewer initial examples leads to better models than trying to curate and train only on more human-curated few-shot examples.

### v2

@@ 4.5.1 Metrics
QA Tasks: We evaluate multiple-choice QA tasks using accuracy, following MMLU’s approach of assessing logarithmic probabilities for vocabulary tokens corresponding to answer labels (Hendrycks et al., 2021). We perform greedy decoding without temperature scaling across answer choices ranging from A-B to A-E.
Generation Tasks: While automated metrics like ROUGE (Lin, 2004) and METEOR (Banerjee and Lavie, 2005) are efficient, their reliance on n-gram overlap limits effectiveness for generative tasks (Barbella et al., 2021). Reference text quality issues and length disparities further reduce reliability (Graham, 2015; Sai et al., 2019; Celikyilmaz et al., 2020). For example, Sottana et al. (2023) found human reviewers often rank benchmark answers among the worst options.
We instead evaluate generations using LLMs as judges (Eldan and Li, 2023), where models provide binary preference scores for output pairs, yielding win rates as metrics (Chiang et al., 2024). This approach shows high inter-rater reliability comparable to human annotators (Hackl et al., 2023; Sottana et al., 2023; Liu et al., 2023).
For general-purpose outputs, we use the popular Alpaca-Eval benchmark (Li et al., 2023) that evaluates multiple LLMs on about 650 human-curated questions (Dubois et al., 2023). We select Llama 3 70B (Dubey et al., 2024) as our annotator model due to its open nature and cost-efficiency for high-volume experiments. As of January 2025, Llama 3 70B ranks 4th in human agreement with a score of 67.5, close to customized GPT-4 versions at 69.2.
@@ 4.5.2 Datasets
We benchmark all models on the original test splits of the human-annotated datasets used for the in-domain (ID) baselines:
BioQA: 397 test samples without images from ScienceQA’s biology subset (Lu et al., 2022).
MedQA: 4,183 validation samples from MedMCQA (Pal et al., 2022).
CSQA: 2,541 validation samples from CommonsenseQA 2.0 (Talmor et al., 2021).
RecipeGen: 1,000 high-quality samples from RecipeNLG (Bień et al., 2020).
Summarization: 1,000 test samples from CNN-DailyMail (See et al., 2017).
5 Results
5.1 Scaling the Data
Figure 3 shows performance improvements from data scaling, reporting means and standard deviations over three random seeds. We observe consistent gains across four tasks with increasing data size relative to the few-shot baseline: 22% (from 65.7 to 80.4), 12% (from 55.3 to 62.1), 23% (from 39.1 to 48.0), and 124% (from 43.7 to 97.9) for BioQA, CSQA, MedQA, and Summarization, respectively, from XS to XL. These results demonstrate CRAFT’s effectiveness across diverse tasks starting from minimal curated examples. Models show appropriate scaling from 100 to 25,000 synthetic samples. Additionally, models trained with fewer examples (32, 100) exhibit higher variance than those trained with 5,000 and 25,000 examples, as indicated by the shaded regions that visualize the standard deviation in the plots.
Notably, CRAFT matches or exceeds Instruct performance across all tasks except RecipeGen. [^1] We further investigate the RecipeGen results in §5.5. It is worth noting that CRAFT uses an LLM in a limited way (to restructure and rewrite existing corpora) that seems to exclude the possibility that distillation has played a role. However, even if distillation were to be considered the reason for good CRAFT performance, the results indicate otherwise: we use the same model as Instruct, Mistral 7B Instruct v0.2, to paraphrase existing corpora examples but achieve even stronger results.
Finally, we observe that CRAFT models outperform the In-Domain (ID) baseline of 25,000 samples in summarization. For other tasks, while we observe lower performance than ID with sample numbers between 1,192 for BioQA, 9,264 for CSQA, and 25,000 for MedQA, we speculate that this could be due to in-domain evaluation for human-curated data. We use their test split to evaluate our models, which may give these models an unfair advantage. We investigate this further in the next section.
5.2 Data Contamination and OOD Generalization
For the experiments in Figure 3, we investigate potential data contamination between test and training examples. We conduct 5-gram weighted Jaccard similarity analyses between CRAFT or in-domain (ID) datasets and the test dataset. For each sample, we combine the instruction and output and gather 5-gram frequencies for the whole dataset. We then calculate the Jaccard similarity between the 5-gram frequency distributions of the respective CRAFT/ID and test datasets, where n-grams receive weight proportional to their frequency.
Dataset In-Domain (ID) CRAFTXL
ScienceQA (In-Domain) 92.7 80.4
MMLUMedical Genetics 67.0 66.0
MMLUAnatomy 58.5 60.0
MMLUHigh School Biology 68.1 69.0
MMLUCollege Biology 69.4 70.8
MMLU-Avg 65.8 66.5
Table 1: Out-of-domain performance of in-domain (ID) vs CRAFTXL models on biology QA tasks. While ID outperforms on its in-domain ScienceQA test set, CRAFTXL shows better generalization to three of four OOD biology subsets from MMLU.
Figure 4: Performance comparison when CRAFT’s retrieval process is initiated with standard human-curated (CRAFT), in-domain (-ID) and purely synthetic (-Synth) few-shots. As the dataset size increases, performance converges across the different few-shot sources, indicating that the retrieval and augmentation framework of CRAFT effectively abstracts away the variability in the quality of the initial few-shots.
This analysis reveals that all CRAFT datasets have less than 0.4% similarity with the test sets, whereas the in-domain datasets show much higher similarities: BioQA (16.6%, 1,192 samples), CSQA (4.4%, 9,264 samples), MedQA (1.1%, 25,000 samples), and Summarization (0.3%, 25,000 samples), indicating some overlap between train and test splits. However, the substantial overlap in in-domain datasets suggests that their reported performance might benefit from train-test similarity.
To isolate generalization capabilities, we evaluate the in-domain (ID) baseline and CRAFT models on four out-of-domain (OOD) biology QA tasks from MMLU (Hendrycks et al., 2021). Table 1 shows that while the ID model outperforms CRAFTXL by 12.3 percentage points on its in-domain test set, CRAFTXL performs better on three of the four OOD biology tasks. This indicates that CRAFT’s training datasets offer better generalization capability and robustness across different domains than human-annotated datasets.
5.3 Sensitivity Analyses
We investigate the sensitivity of the CRAFT pipeline based on different setups: the source of few-shots, the number of few-shots, and the base model used for fine-tuning.
Figure 5: Performance of CRAFT versus Evol-Instruct and Self-Instruct across tasks and dataset sizes with 8 few-shots. CRAFT shows better scaling and higher accuracy than both baselines in most settings.
5.3.1 Few-Shot Source
To assess the sensitivity of CRAFT to the source of the initial few-shots, we evaluate two additional variants alongside our standard human-curated few-shots. In the in-domain (ID) setting, we directly select few-shots from the training splits of our evaluation datasets and manually pair them with appropriate web texts following the design illustrated in Figure 2. This setup represents an optimal distribution match to the test set. The synthetic (Synth) setting employs zero-shot prompts using Mistral 7B Instruct v0.2 to generate both the few-shots and the corresponding web texts without further curation. Although this approach minimizes human effort, it carries the risk of producing repetitive or lower-quality examples.
We run the full CRAFT pipeline using these few-shot variants, fine-tuning Mistral 7B v0.2 and evaluating performance on our standard benchmarks. As shown in Figure 4, most tasks yield very similar performance regardless of the few-shot source. For CSQA and MedQA, performance varies by at most 1.02 and 0.96 percentage points (pp.), respectively, confirming the stability of our approach. Summarization exhibits higher variance across few-shot sources, with a 9 pp. difference favoring manually curated CRAFT few-shots. However, when scaled to CRAFTXL, different few-shot sources lead to more similar results, despite the initially higher variability. Overall, these findings highlight the robustness of CRAFT, demonstrating that even with synthetic few-shots, it consistently generates high-quality, task-specific datasets.
Num FS BioQA CSQA MedQA Summ
8 74,31 60,49 46,83 98,70
32 79,85 61,39 47,50 98,40
Table 2: Accuracy and win rate of CRAFTXL when initiated with 8 and 32 human-curated few-shots. Using just 8 few-shots to initialize CRAFT achieves comparable performance in three of four tasks.
5.3.2 Number of Few-Shots
To better understand the impact of initial human effort required on our synthetic data generation process, we investigate the sensitivity of CRAFT to the number of few-shot examples provided.
Table 2 compares the performance of CRAFTXL, which generates 25,000 synthetic samples, when our retrieval started with 8 versus 32 human-curated few-shots. The results indicate that, while increasing the number of few-shots can yield notable improvements (e.g., BioQA increases from 74.31 to 79.85), the overall performance across three of four tested tasks remains largely comparable. In commonsense QA, medicine QA, and summarization, differences are minimal, with performances varying below 1 percentage point or even matching performance between the two settings. These findings suggest that initializing CRAFT with as few as 8 few-shots is a valid and cost-effective option, significantly reducing the human effort required while still producing high-quality synthetic datasets. Detailed sensitivity results for all dataset sizes (from XS to XL) are provided in Appendix B.
5.3.3 Base Model
In previous sections, we fine-tuned CRAFT models using the pretrained Mistral 7B model. Now, we repeat the experiments using the pretrained Llama 3 8B (Dubey et al., 2024) model. We observe similar trends across all tasks, and the relative improvement is comparable when scaling up from few-shots to 25,000 examples, as illustrated in Table 3. Notably, the datasets created with Mistral 7B Instruct v0.2 are also effective in improving a model with a much stronger baseline performance, such as Llama 3 8B. This finding underscores that our framework leverages the sample-generating LLM only in a reduced capacity – only to augment the retrieved documents – without relying heavily on its inherent performance.
Task CRAFTXS CRAFTS CRAFTM CRAFTL CRAFTXL
BioQA 61.5 64.7 68.8 73.0 78.3
CSQA 55.8 54.8 58.6 60.4 61.4
MedQA 49.5 49.6 52.2 51.5 53.2
Summ 37.3 32.7 86.6 96.8 96.9
Table 3: Accuracy and win rate of CRAFT when Llama 3 8B is used as base model. Datasets generated using Mistral 7B v0.2 can further improve a stronger baseline model such as Llama 3 8B.
5.4 Synthetic Data Generation Comparisons
Figure 5 shows the results when CRAFT is compared against other synthetic data generation methods. CRAFT consistently outperforms Evol-Instruct (EI) across all tasks and dataset sizes. While Self-Instruct (SI) sometimes matches CRAFT’s performance at smaller scales (S/M), CRAFT always achieves higher performance than SI at larger scales (L/XL), except for CSQA at the XL size. Moreover, CRAFT scales more consistently while performing better than the other methods. We believe that a fully distillation-based technique like SI could be competitive with CRAFT in some cases, especially if the number of few-shot examples were increased; however, the available corpus samples in CRAFT improve quality and diversity even with a small number of few-shots.
Importantly, we observed significant differences in method sensitivity across few-shot sources. When running the methods with three different few-shot sources (human-curated, in-domain, and purely synthetic examples), CRAFT demonstrated notably lower variability. At the XL model size, CRAFT achieved an average standard deviation of 1.57 across all five tasks, compared to 3.53 for EI and 4.31 for SI, while also achieving better performance.
5.5 Negative Results: Recipe Generation
Out of our five tasks, we observe non-scaling behavior in one: Recipe Generation. While our manually curated few-shots are of high quality, we see a drop when scaling from 32 to 25,000 examples, as illustrated in Figure 6. CRAFT’s performance is still better than the baseline with official human data, which means that the final dataset is usable. However, we explore why this reverse scaling occurs and examine the drop in performance.
Figure 6: Non-scaling behavior in Recipe Generation: Dataset quality declines with increasing size (100 to 25K examples), showing an inverse scaling trend. This anomaly reflects diminishing data quality at scale.
Generation Method Generated Sample
Human (ID) Sample 1: Select the temperate deciduous forest ecosystem.
A. This ecosystem has: long, cold winters and short, cool summers; many evergreen trees; soil that is poor in nutrients
B. This ecosystem has: year-round rain and warm temperatures; soil that is poor in nutrients; many different types of organisms
C. This ecosystem has: warm, wet summers and cold, wet winters; soil that is rich in nutrients; only a few types of trees  Answer: C
Sample 2: Which organ controls the function of other body organs?
A. skeleton  B. lungs  C. heart  D. brain  Answer: D
CRAFT Sample 1: Which organ plays the primary role in excreting acid and regulating electrolyte balance in the body?
A. Lungs  B. Skin  C. Kidneys  D. Liver  Answer: C
Sample 2: Which simple, single-celled organism formed the first stromatolites on Earth approximately 2 billion years ago?
A. Cyanobacteria  B. Algae  C. Fungi  D. Protozoa  Answer: A
Self-Instruct (SI) Sample 1: Which of the following is a characteristic feature of prokaryotic cells?
A. Nucleus and mitochondria are present in the cell.  B. The cell membrane is not present.
C. The DNA is located in the nucleolus.  D. The cell wall is absent.  Answer: D
Sample 2: Which of the following is a function of the endoplasmic reticulum (ER)?
A. It acts as a site for protein synthesis and folding.  B. It functions in the storage and transport of lipids.
C. It plays a role in the breakdown of complex carbohydrates.  D. It is involved in the process of cell division.  Answer: A
Evol-Instruct (EI) Sample 1: In the early stages of cellular evolution, how did organisms with rudimentary structures develop a complex and dynamic […]
In the pursuit of resource optimization, how did these ancient organisms establish an intricate signaling network, enabling […]
Sample 2: As you delve deeper into the intricate diaphyseal region of a human right femur, you’ll encounter various fascinating […]
Further studies on the presence, distribution, and function of bone marrow […]
Table 4: Qualitative comparison of BioQA samples. For each source, we show two examples to illustrate strengths or shortcomings. CRAFT’s outputs are well-formed and comparable to the human baseline. Self-Instruct maintains the format, but Sample 1 contains a factual error, and Sample 2 includes multiple correct options, making Option B ambiguous in a single-answer context. Evol-Instruct deviates from the task, failing to produce usable QA samples.
An initial analysis suggested that the CRAFT pipeline tends to find less relevant examples over time. We conducted automated data quality analysis to analyze this on a larger scale. For 500 randomly sampled instructions from different sizes of CRAFT datasets (i.e., the training sets), we used Llama 3 8B Instruct to answer the instructions. Then, using Llama 3 70B Instruct as a judge, we compared win rates, i.e., which output the model preferred: the gold output in the CRAFT datasets or the output generated by Llama 3 8B Instruct. We report the average win rate against the Llama outputs as the data quality metric. Higher scores indicate that the pipeline created higher quality output than Llama 3 8B Instruct’s answers.
We observe a decline in data quality when scaling to 25,000 examples: The 100- and 500-example sets achieve win rates of around 0.4, while the 25K set drops to 0.3. This degradation likely causes the performance decline during scaling. We also observe a similar trend in other synthetic data generation methods like Self-Instruct and Evol-Instruct. While the final CRAFT dataset remains practical (outperforming the baseline using official human data), future work should incorporate stopping criteria or more quality validation.
6 Qualitative Analysis
To better understand the quantitative results presented in Section 5, we perform qualitative case studies by inspecting synthesized samples from multiple tasks and data generation methods. This analysis provides intuition as to why certain methods perform better than others and illustrates the style and patterns in generated samples.
6.1 Case Study Design
Our analysis includes both QA and generative tasks across samples generated by CRAFT, Self-Instruct (SI), Evol-Instruct (EI), and human annotators. We use BioQA as the representative for QA tasks, but analyze samples from summarization and recipe generation separately due to format differences. After manually inspecting a wide range of outputs, we selected representative examples that either adhere to or violate the following three criteria: (i) alignment with the expected task format, (ii) content quality and factuality, as well as (iii) stylistic appropriateness. This setup illustrates common advantages and shortcomings.
6.2 Question-Answering Observations
Adherence to format:
As Table 4 shows, both CRAFT and SI produce well-formed QA pairs in a concise format, matching the human in-domain data. EI often drifts into free-form essays, bundles multiple questions into one instruction, and often reaches maximum generation length before completing a sample. The method’s goal to evolve samples by increasing complexity leads to a breakdown in task adherence, which observably resulted in decreased performance in Section 5.
Domain fidelity and correctness:
CRAFT leverages retrieved corpus passages to embed precise terminology (“stromatolites”, “electrolyte balance”) and produces correct QA pairs. Distractor options are semantically proximate yet unambiguously wrong, mirroring expert-curated style. SI’s Sample 1 is factually incorrect, stating that the cell wall is absent in prokaryotes. This exemplifies the risk of ungrounded generation: the model produces plausible-sounding samples at the cost of factual integrity. Furthermore, in SI’s Sample 2, the distractor option B is also correct, creating annotation noise. EI’s questions are verbose and often include multiple sub-questions, making it unclear which question should or will be answered.
6.3 Generative Tasks Observations
Recipe Generation:
The human-annotated in-domain recipes are terse, often just listing steps without providing descriptive details.
Add all dry ingredients together and mix well. Add remaining ingredients stirring only until moistened. […] (Human (ID), Banana Nut Muffin)
In contrast, CRAFT can produce specific, multi-step recipes with headings and detailed directions.
Cook the apples: In a saucepan, cook the sliced apples with sugar over medium heat until softened and caramelized. […]
Add apple filling: Place the apple filling in a line down the middle of the pastry, leaving a 2-inch border on all sides.
Brush the edges: Brush the edges of the pastry with the beaten egg and water mixture.
[…] (CRAFT, Apple Strudel)
CRAFT’s ability to generate more descriptive content stems from grounding the sample synthesis in more elaborate retrieval documents such as cooking blogs. However, as CRAFT retrieves less relevant documents at larger scales, its outputs can degrade. Instructions can become vague (e.g., “Prepare a dish using items from your well-stocked pantry, freezer, and fridge.”) or convoluted, possibly from combining multiple separate recipes from a retrieved document into one (e.g., “Create a salad or bowl by combining greens, proteins, vegetables, cheeses, fruits, and nuts with the specified dressing.”, resulting in an ingredient list that starts with greens but ends with “assorted fruits” and a “scoop of ice cream”. Complete recipe samples are shown in Appendix C.1.
@@ 6.3 Generative Tasks Observations
The human-annotated summaries are written in a headline-driven style, presenting information as extractive bullet points (e.g., “NEW: California Public Utilities Commission passed ban Thursday.”. While factually grounded, this format is limited in stylistic variety and can lack context as well as narrative coherence. CRAFT learns to generate more abstractive, prose-style summaries that are tailored to specific instructional details by being exposed to varying retrieval contexts. The full sample is provided in Appendix C.2.
Instruction: Summarize the text below, focusing on Bill Ackman’s investment in Procter & Gamble and his intentions for the company. Text: [153 word text …]. Summary: Bill Ackman, a hedge fund manager, believes that Procter & Gamble’s (P&G) stock could be worth more than its current price due to poor marketing and pricing strategies. He holds around 1% of the company’s shares and is actively pushing for changes. [48 more words …]
SI also demonstrates task adherence and is likewise able to adjust its style toward varying instructions (e.g., summarizing a text “in a tweet”), but its outputs tend to be generic and lack the specificity seen in CRAFT’s document-grounded summaries. EI again introduces superficially complex phrasing, such as “Unraveling the Perplexing Enigma of Age-Related Pituitary […]”, or “In the intricately woven narrative of Afghanistan’s rich and complex history […]”.